### v2

# Poker Hand History File Format Specification

Abstract
This paper introduces the Poker Hand History (PHH) file format, designed to standardize the recording of poker hands across different game variants. Despite poker’s widespread popularity in the mainstream culture as a mind sport and its prominence in the field of artificial intelligence (AI) research as a benchmark for imperfect information AI agents, it lacks a consistent format that humans can use to document poker hands across different variants that can also easily be parsed by machines. To address this gap in the literature, we propose the PHH format which provides a concise human-readable machine-friendly representation of hand history that comprehensively captures various details of the hand, ranging from initial game parameters and actions to contextual parameters including but not limited to the venue, players, and time control information. In the supplementary, we provide over 10,000 hands covering 11 different variants in the PHH format. Building on our previous work on PokerKit, a premier poker hand simulation tool, we demonstrate the usages of our open-source Python implementation of the PHH parser. The source code of the parser is available on GitHub: https://github.com/uoftcprg/pokerkit
Index Terms:
Card games, Games of chance, Game design, Multi-agent systems, Poker, Rule based systems, Scripting, Strategy games
I Introduction
Poker is a popular card game with a stochastic nature and imperfect information aspect that has not only captured the mainstream attention as a leading mind sport but also served as a benchmark of artificial intelligence (AI) research, namely in the development of imperfect information AI agents.
Pioneering breakthroughs in imperfect information game-solving algorithms have often been accompanied by their applications in state-of-the-art poker AI agents as seen in the example of Rhode Island hold ’em [1], fixed-limit Texas hold ’em [2], heads-up no-limit Texas hold ’em [3, 4], and 6-player no-limit Texas hold ’em [5]. Successive state-of-the-art algorithms were shown to solve games of progressively larger game tree sizes.
While impressive, the applications of these algorithms were mostly limited to the Texas hold ’em variant. There exists countless variants of poker that introduce different rules and actions not present in Texas hold ’em [6, 7] that can serve as new challenges in the development of imperfect information AI agents. For instance, Omaha variants involve four hole cards per player instead of simply two, which increases the range of possible holdings of opponents, further increasing the game tree size. Another popular variant, lowball draw, introduces shifting hidden information depending on the discards and draws unlike Texas hold ’em where the players’ hole cards remain unchanged throughout the hand.
The lack of exploration in the aforementioned areas of computer poker can partly be explained by the lack of reliable open-source multivariant poker tooling. Our recent work on PokerKit [8] aimed to provide a robust implementation of poker simulation and hand evaluation suite to address this absence and has since established itself as a popular tool in computational poker.
However, our previous efforts in the digitalization of poker were only partially complete, as we provided no means to record poker hands for persistent storage and subsequent analysis. This is partly due to the fact that there does not exist a standardized format for recording poker games that is not only human-friendly but also easily parsable by computer software akin to the portable game notation (PGN) for chess [9] or smart game format (SGF) for go [10].
This paper proposes a novel open Poker Hand History (PHH) file format designed to provide a consistent method to annotate poker hands across different variants in addition to capturing other details of the game including but not limited to the venue, players, and time control information that can aid in domains such as poker AI agent development, historical analysis, and data referencing.
In the next section, we provide some motivational examples that showcase the versatility of the PHH notation. This is followed by an exploration of the related works in the literature that introduce alternatives to our proposal with which we later run benchmarks for comparison purposes. Then, we provide the detailed specification of the PHH format, listing the required and optional fields used to comprehensively describe poker hands in addition to various recommendations in file formatting, style guides, and parser implementation. Finally, for demonstrative purposes, we provide sample usages of our open-source Python PHH reader and writer which uses PokerKit as the simulation backend.
In the supplementary, over 10,000 sample poker hands, covering 11 different poker variants and composed of small selections of historical or notable hands, all 83 televised hands in the 2023 WSOP Event #43: \$50,000 Poker Players Championship (PPC) — Day 5 [11], and the hands played by Pluribus in Brown and Sandholm [5], are provided in the PHH file format.
II Motivational Examples
# A big pot between Dwan and Ivey.
variant = "NT"
antes = [500, 500, 500]
blinds_or_straddles = [1000, 2000, 0]
min_bet = 2000
starting_stacks = [
  1125600,
  2000000,
  553500,
actions = [
  # Pre-flop
  "d dh p1 Ac2d",  # Ivey
  "d dh p2 ????",  # Antonius
  "d dh p3 7h6h",  # Dwan
  "p3 cbr 7000",  # Dwan
  "p1 cbr 23000",  # Ivey
  "p2 f",  # Antonius
  "p3 cc",  # Dwan
  # Flop
  "d db Jc3d5c",
  "p1 cbr 35000",  # Ivey
  "p3 cc",  # Dwan
  # Turn
  "d db 4h",
  "p1 cbr 90000",  # Ivey
  "p3 cbr 232600",  # Dwan
  "p1 cbr 1067100",  # Ivey
  "p3 cc",  # Dwan
  "p1 sm Ac2d",  # Ivey
  "p3 sm 7h6h",  # Dwan
  # River
  "d db Jh",
author = "Juho Kim"
event = "Million Dollar Cash Game"
year = 2009
players = [
  "Phil Ivey",
  "Patrik Antonius",
  "Tom Dwan",
currency = "USD"
Figure 1: The first televised million-dollar pot between Tom Dwan and Phil Ivey [12] in the PHH file format. Note that the starting stack of Patrik Antonius is unknown.
# A bad beat between Yockey and Arieh.
variant = "F2L3D"
antes = [0, 0, 0, 0]
blinds_or_straddles = [
  75000, 150000, 0, 0,
small_bet = 150000
big_bet = 300000
starting_stacks = [
  1180000, 4340000, 5910000, 10765000,
actions = [
  "d dh p1 7h6c4c3d2c",  # Yockey
  "d dh p2 ??????????",  # Hui
  "d dh p3 ??????????",  # Esposito
  "d dh p4 AsQs6s5c3c",  # Arieh
  "p3 f",  # Esposito
  "p4 cbr 300000",  # Arieh
  "p1 cbr 450000",  # Yockey
  "p2 f",  # Hui
  "p4 cc",  # Arieh
  "p1 sd",  # First draw; Yockey
  "p4 sd AsQs",  # Arieh
  "d dh p4 2hQh",  # Arieh
  "p1 cbr 150000",  # Yockey
  "p4 cc",  # Arieh
  "p1 sd",  # Second draw; Yockey
  "p4 sd Qh",  # Arieh
  "d dh p4 4d",  # Arieh
  "p1 cbr 300000",  # Yockey
  "p4 cc",  # Arieh
  "p1 sd",  # Third draw; Yockey
  "p4 sd 6s",  # Arieh
  "d dh p4 7c",  # Arieh
  "p1 cbr 280000",  # Yockey
  "p4 cc",  # Arieh
  "p1 sm 7h6c4c3d2c",  # Showdown; Yockey
  "p4 sm 2h4d7c5c3c",  # Arieh
author = "Juho Kim"
event = "2019 WSOP Event #58"
city = "Las Vegas"
region = "Nevada"
country = "United States of America"
day = 28
month = 6
year = 2019
players = [
  "Bryce Yockey", "Phil Hui",
  "John Esposito", "Josh Arieh",
Figure 2: A bad beat between Josh Arieh and Bryce Yockey [13] in the PHH file format.
Figures 1 and 2 show two historical poker hands digitalized into the PHH format. The former involves the first televised million-dollar pot between Tom Dwan and Phil Ivey in the 2009 Full Tilt Million Dollar Cash Game Season 4 Episode 12 [12]. The latter describes a bad beat from the final table of the 2019 WSOP Event #52: \$50,000 PPC on Day 5 of the event between Josh Arieh and Bryce Yockey [13].
The PHH format aims to be both human-readable and writable while maintaining easy parsability by relevant software systems. In this format, the hands are described through the definitions of various key/values, along with optional comments throughout the file.
III Related Works
In poker literature, it has historically been the norm to describe poker games verbally as exemplified in the news articles covering the aforementioned hands [12, 13]. Such a practice, while fit for human readership, presents significant challenges for parsing and computational analysis. While the description could be rich with information like the names of players involved, the backstory of the hand, et cetera, it could also lack nuances and details that may be of interest to the readers such as the exact stack sizes, bet sizes, and mucked hand values, just to name a few. The subjective formatting may introduce bias and inconsistencies unsuitable for scientific use cases.
One of the earliest attempts to introduce a form of structure to hand summaries was the Mike Caro University Poker Chart proposed by Mike Caro [14]. His tabular format where players and betting rounds are represented as a table adds a form of structure to the data that makes it easier for humans to follow. In addition, various mathematical symbols and operations such as “=,” “$\sqrt{}$,” “–,” exponentiation, and more were used to concisely represent each betting action. Caro’s format serves as an innovative breakthrough when it comes to improvements in visual clarity and consistency. Unfortunately, various descriptions of poker such as stakes remain verbal, and are designed primarily for human-to-human communication. In hindsight, it failed to become a standard in the poker community, especially in the age of computer poker where machine interpretability is another vital angle that must be considered when designing such standards.
PokerStars Hand #????????????: Holdem No Limit (\$1000.00/\$2000.00) - 2009/??/?? ??:??:?? ??? [2009/??/?? ??:??:?? ET]
Table ‘?????’ ?-max Seat #3 is the button
Seat 1: PhilIvey (\$1125600.00 in chips)
Seat 2: PatrikAntonius (\$2000000.00 in chips)
Seat 3: Hero (\$553500.00 in chips)
PhilIvey: posts the ante \$500.00
PatrikAntonius: posts the ante \$500.00
Hero: posts the ante \$500.00
PhilIvey: posts small blind \$1000.00
PatrikAntonius: posts big blind \$2000.00
*** HOLE CARDS ***
Dealt to Hero [7h 6h]
Hero: raises \$5000.00 to \$7000.00
PhilIvey: raises \$16000.00 to \$23000.00
PatrikAntonius: folds
Hero: calls \$16000.00
*** FLOP *** [Jc 3d 5c]
PhilIvey: bets \$35000.00
Hero: calls \$35000.00
*** TURN *** [Jc 3d 5c] [4h]
PhilIvey: bets \$90000.00
Hero: raises \$142600.00 to \$232600.00
PhilIvey: raises \$834500.00 to \$1067100.00 and is all-in
Hero: calls \$262400.00
Uncalled bet (\$572100.00) returned to PhilIvey
*** RIVER *** [Jc 3d 5c 4h] [Jh]
*** SHOW DOWN ***
PhilIvey: shows [Ac 2d] (a straight, Ace to Five)
Hero: shows [7h 6h] (a straight, Three to Seven)
Hero collected \$1109500.00 from pot
*** SUMMARY ***
Total pot \$1109500.00 | Rake ?
Board [Jc 3d 5c 4h Jh]
Figure 3: The identical hand played in Figure 1 converted to the PokerStars hand history format. Note that unknown information is marked with question marks and the first line was broken to fit in the page. Again, the starting stack of Patrik Antonius is unknown.
PokerStars, one of the most popular online poker platforms, saves a log of the game history in their PokerStars Hand History format [15]. As an example, the identical hand shown in Figure 1, converted into the PokerStars hand history format, is shown in Figure 3. While the logs are designed for human readership and record-keeping, they are in plaintext format and are highly structured and consistent, allowing various poker analytic tools in the industry such as Holdem Manager and Poker Copilot to parse and analyze a player’s hand history reliably. The logged details include granular information down to the amounts of antes and blinds being posted by each player. This fine detail unfortunately makes it too verbose to be hand-typed by a human while being too specialized for use in video poker. For instance, some digital archiving use cases may require additional or complete descriptions of the game like the final stacks, time control, et cetera. Not only that, the verbal nature of the format introduces unnecessary verbosity that wastes storage space.
STATE:7:fr225fffc/cr475c/cr1225c/cc:5hJc|Jd9h|6s5c|Ah7h|2s2d|3hTs/3sJh2h/Tc/Ks:-50|1275|0|-1225|0|0:Gogo|Budd|Eddie|Bill|Pluribus|MrWhite
Figure 4: A sample hand played by Pluribus in the supplementary of Brown and Sandholm [5]. Note that the line was broken to fit in the page.
The dataset of hands played by Pluribus, in the supplementary of Brown and Sandholm [5], is formatted in a much more concise way as shown in Figure 4. This format includes the bare minimum description of actions to comprehensively represent a no-limit Texas hold ’em hand. However, due to the fact that each action notation does not specify the actor, the hands are difficult to follow by a human reader. In addition, the format is limited to only a single variant: no-limit Texas hold ’em, and makes critical assumptions such as the identical and fixed starting stack sizes of 100 big blinds and the number of players being 6. On top of this, it lacks options to include various metadata outside of the game state descriptions that may be of interest for data referencing purposes.
Standardized recording formats of other prominent board games such as PGN format for chess [9] and SGF for go [10] provide a good direction that can be followed for the game of poker. These formats specify tags or properties that can be omitted depending on the needs of the end user for encoding information not limited to the game state or the list of actions but also numerous extra information designed to give extra context to the game played like the location or the commentary just to name a few. Our PHH format takes inspiration from the aforementioned approaches and adopts them for use in the recording of poker hand history.
IV Objective
An ideal representation of poker hands should not only be concise but also be easily readable and writable by human readers. It should also support only describing the bare minimum information necessary to comprehensively describe a hand while facilitating options to give rich details for various use cases beyond simple action tracking. For convenience in parser implementations, the formatting must also be simply structured in a consistent manner. In addition, the format must be powerful enough to describe a wide variety of poker games played in formal settings. Such selections should cover games played in reputable poker tournaments around the world while the inclusion of various eccentric or esoteric variants is not as necessary.
V Specification
The PHH format is a derivative of the Tom’s Obvious, Minimal Language (TOML) format [16]. As such, all PHH files must also be valid TOML files, following the TOML syntax. This design allows PHH files to be human-readable and writable while maintaining easy parsability by software systems. Software parser implementations are encouraged to utilize built-in or third-party tools to read and write TOML files.
The PHH format puts restrictions on the naming and types of the key/values, with which the poker games are described. These definitions, named “fields,” fall under the following three classifications:
  1. 1.
    Initial state construction.
  2. 2.
    Subsequent state progression.
  3. 3.
    Miscellaneous game information.
The first and second classifications, involving state construction and progression, are grouped together as “required” fields and must be specified. The third classification is dubbed as “optional” fields and may be omitted by the annotator.
V-A Error Handling
In the implementation of the PHH parser, the program should report an error when any violation of the specification is encountered. The parser may stop parsing the file, ignore the problematic field or action, or proceed as normal.
V-B Position
The fields that have a value of type array where each element corresponds to a particular player must have their elements ordered identically to the ordering of the players and have lengths equal to the number of players. The ordering of the players is table agnostic, that is, it is purely based on position.
The first player in the PHH format is typically the very first person being dealt the hole card by the dealer while the last player in the PHH format is typically the person being dealt the last hole card. This is assuming a proper dealing order is followed. In general, the ordering of the players must represent a clockwise order.
In typical button games such as no-limit Texas hold ’em or pot-limit Omaha hold ’em, the first player must be in the small blind while the last player must be in position (i.e. has the button). In non-button games like stud, there is no strict notion of position. In such games, the first player should be the one in the dealer’s immediate left while the last player should be in the dealer’s immediate right.
V-C Rules
As poker is a game with a diverse set of rules that often differ based on regions or events, it is impossible for us to point to anything as an authoritative documentation of poker rules that must be followed both by the annotator and the parser. The parser implementation should be designed to handle actions that may be potential transgressions in certain poker rule sets, either silently or with a warning.
While annotating the fields, especially the “actions” field, the author should follow the specific set of rules followed by the game being annotated. Where applicable or possible, the author is encouraged to conform to the latest WSOP Tournament Rules and WSOP Live Action Rules of which the most recently released versions as of the writing of this paper are for the 2023 WSOP [6, 7]. If there is any contradiction between the two documents, the tournament rules take precedence, as they impose stricter guidelines on poker playing. When there is an ambiguity in both documents, it is to be considered as an undefined behavior. In such a case, the annotator should adhere to a historically informed custom of poker playing.
VI Required Fields
TABLE I: The 11 variant codes and the corresponding full variant names.
Code Name
“FT” Fixed-limit Texas hold ’em
“NT” No-limit Texas hold ’em
“NS” No-limit short-deck hold ’em
“PO” Pot-limit Omaha hold ’em
“FO/8” Fixed-limit Omaha hold ’em high/low-split eight or better
“F7S” Fixed-limit seven card stud
“F7S/8” Fixed-limit seven card stud high/low-split eight or better
“FR” Fixed-limit razz
“N2L1D” No-limit deuce-to-seven lowball single draw
“F2L3D” Fixed-limit deuce-to-seven lowball triple draw
“FB” Fixed-limit badugi
TABLE II: The composition of required fields for each variant.
Variant Antes Blinds or Straddles Bring-in Small bet Big bet Min bet Starting stacks Actions
“FT” yes yes no yes yes no yes yes
“NT” yes yes no no no yes yes yes
“NS” yes yes no no no yes yes yes
“PO” yes yes no no no yes yes yes
“FO/8” yes yes no yes yes no yes yes
“F7S” yes no yes yes yes no yes yes
“F7S/8” yes no yes yes yes no yes yes
“FR” yes no yes yes yes no yes yes
“N2L1D” yes yes no no no yes yes yes
“F2L3D” yes yes no yes yes no yes yes
“FB” yes yes no yes yes no yes yes
TABLE III: The required fields and their TOML native types.
Field Name TOML Native Type
Variant variant String
Antes antes Array of integers or floats
Blinds or Straddles blinds_or_straddles Array of integers or floats
Bring-in bring_in Integer or float
Small bet small_bet Integer or float
Big bet big_bet Integer or float
Min bet min_bet Integer or float
Starting stacks starting_stacks Array of integers or floats
Actions actions Array of strings
The required fields represent a handful of fields that the annotator must define to digitalize a poker hand and therefore form a valid PHH file. The required fields are either used during the initial state construction or during the subsequent state progression.
In this paper and the supplementary, 11 variants are explored, each of which is written in an abbreviated form, as tabulated in Table I. These variants include all 9 played in the 2023 WSOP Event 43: \$50,000 PPC [11], in addition to 2 others like short-deck and badugi in which the poker community has shown growing interest. These were chosen as they adequately represent the diverse nature of poker variants and demonstrate the flexibility of the specification laid out in this paper which is powerful enough to accommodate all variants in the 2023 World Series of Poker Tournament Rules [6]. Those not mentioned in Table I can simply be integrated by introducing new variant codes to be associated with, and they will be introduced on a need-by-need basis based on community feedback in the future version of the specification.
The actual composition of the required fields varies, depending on the played variant. The required field statuses for each variant are described in Table II, and the fields’ native TOML types are specified in Table III.
VI-A State Construction
The fields in this section concern itself with the topic of the initial state construction.
VI-A1 Variant
The “variant” field, which has a string value, defines what variant the hand is in and therefore determines the composition of the required fields that must be specified. The field value is a code unique to each variant, 11 of which are tabulated in Table I, along with the corresponding field requirements. Sample definitions are shown in Figures 1 and 2.
VI-A2 Antes
The value of the “antes” field, representing the antes, must be an array of non-negative integers or floats of length equal to the number of players. In general, there are two types of antes: uniform antes and non-uniform antes. In the case of uniform antes, all the array elements are identical, meaning the antes are the same for each player. No antes (zero antes) must be represented as an array of zero integers or floats while non-zero uniform antes can be represented as an array of constant positive TOML native numerical values. Non-uniform ante formats such as big-blind antes and button antes must be represented as an array with non-zero values in the indices corresponding to the anted players. Example definitions are shown in Figure 5.
# Zero antes.
antes = [0, 0, 0, 0, 0]
# Antes of 1.
antes = [1, 1]
# Big blind ante
antes = [0.0, 3.0, 0.0]
# Button ante
antes = [0, 0, 0, 10]
Figure 5: Example ante definitions in PHH format.
VI-A3 Blinds or Straddles
The “blinds_or_straddles” field, representing all forced bets except the bring-in, must only be specified in button games as an array of non-negative integers or floats of length equal to the number of players. Most button games have two blinds: a small blind and a big blind. These are simply described by an array with two positive values in the first two indices while all other elements are zero. Optionally, various straddle regimes ranging from under-the-gun straddle, double straddle, button straddle, et cetera, can be introduced by setting the elements in the relevant indices with non-zero values. Examples are demonstrated in Figure 6. Note that, in heads-up situations, the effective blind amounts are the reversed version of the array field value.
# Small and big blind (heads-up)
blinds_or_straddles = [1, 2]
# Small and big blind
blinds_or_straddles = [1, 2, 0, 0, 0, 0]
# Under-the-gun (UTG) straddle
blinds_or_straddles = [1, 2, 4, 0]
# UTG and UTG+1 straddle
blinds_or_straddles = [1, 2, 4, 8, 0, 0]
# Button straddle
blinds_or_straddles = [1, 2, 0, 0, 0, 4]
Figure 6: Example blind or straddle definitions in PHH format.
VI-A4 Bring-in
The “bring_in” field is of a positive integer or float value and is primarily used in stud poker variants to represent the bring-in forced bet. The usage of bring-ins is mutually exclusive with blinds or straddles. In other words, both must never be defined together. A sample definition is shown in Figure 7.
bring_in = 1
small_bet = 2
big_bet = 4
Figure 7: Selections of required field definition in PHH format.
VI-A5 Small bet
The “small_bet” field, a positive integer or a float, denotes the minimum bet in the first few betting rounds of fixed-limit games. It is not a feature of pot-limit or no-limit games. Sample use cases are shown in Figures 2 and 7.
VI-A6 Big bet
The “big_bet” field, a positive integer or a float, similar to its smaller counterpart, denotes the minimum bet in the last few betting rounds of fixed-limit games. Again, this field is not featured in unfixed-limit games. sample definitions are shown in Figures 2 and 7.
VI-A7 Min bet
The “min_bet” field, a positive integer or a float, denotes the minimum bet in all of the betting rounds of no-limit or pot-limit games. This field must not be specified in fixed-limit games. A sample usage is shown in Figure 1.
VI-A8 Starting stacks
The “starting_stacks” field denotes the starting stacks of each player and is an array of positive integers or floats of length equal to the number of players. All stack sizes must strictly be non-zero. Sample usages are shown in Figures 1 and 2.
VI-B State Progression
The one and only field that represents the progression of a poker state is the “actions” field. This field is an array of string action representations, following our action notation described below. It describes how the state mutates throughout the hand as the actors – the dealer and the players – act.
VI-B1 Actions
The “actions” field is an array of string action notations that describe what each actor did in their turn. For the dealer (an actor), their action notations denote what cards were dealt to each player or to the board. More specifically:
  • Dealing board cards or
  • Dealing hole cards.
For each player, their action notations may describe them:
  • Posting a bring-in;
  • Completing;
  • Betting;
  • Raising or raising to;
  • Checking;
  • Calling;
  • Folding;
  • Standing pat;
  • Discarding;
  • Showing their hole cards; or
  • Mucking their hole cards.
The notations in the “actions” field must be sorted in the order of the action. Its first element, if the array is not empty, must be the first action taken, which, in all mainstream poker variants, is the hole card dealing action by the dealer. The actions may represent a complete history of the hand or a partial history that does not reach the terminal state (the hand being over).
Figures 1 and 2, as well as the PHH files in the supplementary, demonstrate the sample usages of this field.
VI-C Action Notation
Each action notation aims to concisely describe the actor, the performed action, and arguments if any. The notation may also optionally include a verbal commentary. In addition, a notation may be composed entirely of a standalone commentary or be an empty string. In general, the action notation satisfies the following grammar:
$[Actor\ Action[\ Arguments...]][\ \#\ Commentary]$
where contents that may be omitted are enclosed in square brackets. The words are separated by single or multiple consecutive whitespace characters, and the notation itself may contain leading or trailing whitespaces.
Note that commentaries and TOML comments are two distinct mechanisms in the PHH format. They are both followed by a single hash symbol. However, TOML comments are ignored by the parser, while the commentary information is considered a part of the action notation by the parsing system. The annotator should use commentaries to give extra contextual information about the hand that cannot be determined by the parser such as a summary of the metagame between some of the players, exposed hole cards, and interruptions, just to name a few, whereas the TOML comments should serve as a guide to the reader who is directly reading the file without the help of a software parser. For instance, the information provided by TOML comments may include the names of the betting round or the actor.
[…]
VI-C1 Board Dealing
[…]
is the board dealing action where the dealer deals the community cards to the board at the beginning of each applicable street. The notation of the board dealing action must include a single argument of a valid card notation representing the newly dealt community cards. The correct number of cards must be dealt; the cards must all be notated as known cards. Sample action notations are written in Figure 1.
VI-C2 Hole Dealing
The hole dealing action
[…]
represents the action of dealing hole cards to a particular player and must accept two arguments, the player being dealt and the newly dealt hole cards (in card notation) some or all of which may be unknown. Its usages are shown in Figures 1 and 2. As in the board dealing action, the correct number of hole cards must be dealt to each player. The hole dealing must be made starting from the first active player to the last active player still in the hand.
VI-C3 Standing Pat or Discarding
The standing pat and discarding actions
[…]
are grouped together due to their similarity. Standing pat is a special case of discarding where a player does not discard any hole cards. This notation accepts an optional argument denoting the cards that are being discarded. An omission of this argument denotes that the player in turn is standing pat while the inclusion denotes that the player discarding. All discarded cards must be part of the player’s hole cards. Therefore, the card notation in the argument may contain unknown cards if and only if some of the player’s hole cards are unknown. This action is demonstrated in Figure 2.
VI-C4 Bring-in Posting
This action, with the
$\text{p}n\text{ pb}[\text{ \# Commentary}]$
grammar, denotes the posting of the bring-in forced bet. A sample notation of such action is: “p3 pb” which describes that Player 3 is bringing-in.
VI-C5 Folding
$\text{p}n\text{ f}[\text{ \# Commentary}]$
is the folding action. Examples are shown in Figures 1 and 2.
VI-C6 Checking or Calling
This action represents checking or calling
$\text{p}n\text{ cc}[\text{ \# Commentary}]$
which entails matching the largest bet. When the amount required to match is zero, the player is said to check. Otherwise (in the case of a non-zero amount), the player is said to call. The examples are shown in Figures 1 and 2.
VI-C7 Completion, Betting, or Raising to
This action represents completion, betting, or raising to a positive integral or real amount, supplied as an argument, and uses the
$\text{p}n\text{ cbr amount}[\text{ \# Commentary}]$
grammar. This positive amount must be between the valid ranges set by the betting structure of the variant. When this action is performed as an alternative to or in the face of a bring-in, the player is said to complete. Alternatively, when nobody else bet, the player is said to be betting. Otherwise, the player is said to be making a raise. These are demonstrated in Figures 1 and 2.
VI-C8 Showing or Mucking Hole Cards
[…]
action is the only action that can be performed during showdown or all-in situations. It represents showing of the hole cards to try to win the pot or after an all-in, or mucking of the hole cards. This action accepts an optional argument of cards which, if supplied, must all be known. The omission of this field represents mucking while the inclusion represents showing. Instances of this action are shown in Figures 1 and 2.
VI-C9 No operations
No-ops
[…]
are represented with an empty string, a string composed entirely of whitespace characters like “ ”, or a standalone commentary like “# Burn card 6s is exposed.”
VI-D Card Notation
TABLE IV: All ranks and their single character representations.
Rank Character
Deuce 2
Trey 3
Four 4
Five 5
Six 6
Seven 7
Eight 8
Nine 9
Ten T
Jack J
Queen Q
King K
Ace A
Unknown ?
TABLE V: All suits and their single character representations.
Suit Character
Club c
Diamond d
Heart h
Spade s
Unknown ?
Any card in the PHH format is either known or unknown. Each known card must be represented with two characters. The first character represents the rank of the card, and must either be one of the numeric characters from “2” to “9”, representing the analogous ranks from deuces to nines, or an uppercase alphabet character “T”, “J”, “Q”, “K”, and “A” for the ranks of tens, jacks, queens, kings, and aces, respectively. The second character represents the suit, the valid values of which must be the lowercase alphabet characters “c”, “d”, “h”, and “s” for clubs, diamonds, hearts, and spades, respectively. All ranks and suits are tabulated in Table IV and V, respectively. An unknown card must be represented with two question mark characters: “??”. Multiple cards must be represented by concatenating individual single-card representation without any separators or delimiters such as “AdAcKh” for ace of diamonds, ace of clubs, and king of hearts.
Certain variants like short-deck hold ’em use a non-standard deck where low ranks from deuces to fives do not exist. When a card that does not exist for a particular variant being played is specified, the parser should treat this as a violation of the specification.
VII Optional Fields
TABLE VI: The optional fields and their TOML native types.
Field Name TOML Native Type
Author author String
Event event String
URL url String
Address address String
City city String
Region region String
Postal code postal_code String
Country country String
Time time Local time
Time zone time_zone String
Day day Integer
Month month Integer
Year year Integer
Hand hand Integer
Level level Integer
Seats seats Array of integers
Number of seats seat_count Integer
Table table Integer
Players players Array of strings
Finishing stacks finishing_stacks Array of integers or floats
Currency currency String
Ante trimming status ante_trimming_status Boolean
Time limit time_limit Integer or float
Time banks time_banks Array of integers or floats
currency = # error
Figure 8: An example syntax error in PHH and TOML files [16].
The optional fields may be left unspecified by the annotator. The complete list of optional fields and their TOML native types are tabulated in Table VI. When a field is included, it must be completely defined. For instance, a partial definition as in Figure 8 is strictly forbidden as it is syntactically incorrect as per the TOML specification [16].
VII-A Annotator
VII-A1 Author
The “author” field describes the annotator and is of TOML native string type. The value should represent a properly capitalized full name of the annotator as in
``$First\_name[\ Middle\_name(s)...]\ Last\_name"$
or a mononym
``$Name"$
depending on their preferences.
VII-B Venue
The following fields describe the venue in which the hand was played.
VII-B1 Event
The “event” field represents the full name of the event which should be in English and is of type: string.
VII-B2 URL
The “url” field contains a string URL relevant to the event. This field is particularly important for online poker hands where the venue might not be a physical location. Still, the annotator may choose to specify a relevant URL to the organizer’s website or tournament website for physical poker hands, as well. The link should be prefixed with substrings like “http://” or “https://” to denote the internet communication protocol.
VII-B3 Address
The “address” field contains a string value that may or may not include the street name, street number, room number, et cetera. The annotator should use English wherever possible.
VII-B4 City
The “city” field contains the information about the city the hand was played in. The annotator should use full English names for the cities.
VII-B5 Region
The “region” field contains the information about the city the hand was played in. Similarly to the cities, the annotator should use full English names for the regions.
VII-B6 Postal Code
The “postal_code” field contains the postal code of the location of the hand.
VII-B7 Country
The “country” field contains the name of the country of the hand location. Again, the annotator should use full English names for the countries.
VII-B8 Time
The “time” field should contain the approximate time the hand began as a TOML native local time without the timezone information.
If the “time_zone” field is defined, the time is assumed to be of that time zone. Otherwise, if every one of the following three fields – “city,” “region,” and “country” – are defined, it is assumed that the time is in the local time zone of the location of the hand. If the two previous conditions are not satisfied, the time is assumed to be in Coordinated Universal Time (UTC).
VII-B9 Time Zone
The “time_zone” field contains a value of type string and must represent an IANA time zone in the latest version of the IANA time zone database [17] available like “America/Toronto.”
VII-B10 Day
The “day” field should contain the day the hand took place in, and is of the integer type.
VII-B11 Month
The “month” field should contain the month the hand took place in, and has an integral value where 1 represents January, 2 represents February, and so on.
VII-B12 Year
The “year” field should contain the year the hand took place in, and must be an integer.
VII-C Counter
VII-C1 Hand
The “hand” field denotes the hand number which should be an integer.
VII-C2 Level
The “level” field denotes the blind or ante level, which is relevant in a poker tournament. This field must contain an integer value.
VII-D Seating
VII-D1 Seats
The “seats” field denotes the seat numbers of the players. The value of this field must be an array of integers, of length $n$ where $n$ is the number of players.
VII-D2 Number of Seats
The “seat_count” field denotes the number of seats and must have an integer value. Note that it is possible for there to be more seats than the players.
VII-D3 Table
The “table” field denotes the table number the hand is played in, as an integer value.
VII-E Miscellaneous
VII-E1 Players
The “players” field contains player names as an array of strings. The names should be written as a full name (“First_name[ Middle_name(s)…] Last_name”) or as a mononym (“Name”) depending on which is used or preferable. If the name of the player is unavailable perhaps due to anonymity, an empty string should be used.
VII-E2 Finishing Stacks
The “finishing_stacks” field denotes the final stacks as an array of non-negative integers or floats. It may also be specified for a non-terminal hand history file, which can be interpreted as the stack values after all the action notations are applied.
The description of finishing stacks is helpful as the parser may not be aware of the granularity of the currency the chips are in. For example, dollars can be broken down to cents whereas Japanese Yen must be of an integral value. On top of this, in a physical setting where chips are used, depending on the denominations, odd chip situations may arise where the player out of position is given the extra odd chip that cannot be broken further. It is, of course, infeasible to describe all the different chip values each player has in a poker hand history format. These inaccuracies are inherent drawbacks of using purely numerical representations to describe the stack values. It is worth noting that the inconsistencies caused by such circumstances only lead to extremely minor ambiguities in the final stack sizes that should not significantly impact the expected value calculations.
VII-E3 Currency
The “currency” field denotes what currency the chips are in. The value must be of string and be one of three letter currency values in ISO 4127 standard [18].
VII-E4 Ante Trimming Status
The “ante_trimming_status” denotes how to handle the special cases where a player or players are so short-stacked that they cannot even pay the full ante as a Boolean value that defaults to “false.” If “true,” the player can only win depending on how much the player contributed. Otherwise, the player can win all the antes even if they did not pay the full ante. This field was introduced due to the ambiguities in the WSOP rules [6, 7]. Note that, in certain ante configurations such as big blind antes, unequal contribution in the antes is expected, and therefore this value should be kept as “false” for non-uniform antes. In the vast majority of real-life poker hands, the players can at least pay the full antes. Therefore, this field is only relevant in extremely rare circumstances.
VII-F Time Control
VII-F1 Time Limit
The “time_limit” represents the shot clock time and is of type: integer or float. It represents the time the user has to make a decision at every turn. If there is none, the annotator can omit this field or set it as “inf” for infinity.
VII-F2 Time Banks
The “time_banks” field denotes the time banks of each player at the beginning of the hand, as an array of integers or floats of length equal to the number of players. Just like in stack values, representing time banks as a simple numerical value loses the detail of the granularity of the time cards a player may have. Still, integral values are used thanks to their simplicity and consistency.
VIII User-defined Fields
_sponsors = ["", "GGPoker", ""]
_nationalities = [
  "Peru", "Italy", "Japan",
Figure 9: Example user-defined fields in PHH format.
Depending on the use case, the required and optional fields may not be sufficient to capture certain desired levels of details, as demonstrated in Figure 9. The specification allows annotators to define custom fields. Parser implementations should simply ignore unknown fields and may issue a warning when such a field is encountered. This is also in part to ensure that the parser is compatible with PHH files conforming to the future version of the specification that may introduce new fields.
The names of the user-defined fields must not be one of the names of the required or optional fields and should be chosen carefully such that it does not conflict with future versions of the specification. The naming conventions of user-defined fields are explained in the style guide.
IX Style Guide
In PHH files, the fields may be in any order, and the formatting may follow any style, as long as it conforms to the TOML syntax [16]. However, the annotator is encouraged to list all required fields first before listing the optional fields and enclose the equal signs “=” for each key/value assignment with a single space character “ ” on both sides to enhance readability.
When naming the user-defined fields, users should use the “snake_case” naming convention where words are written in lowercase characters, separated by an underscore as in the required and optional fields to maintain consistency. In addition, the names of the user-defined fields should start with a single underscore as shown in Figure 9 to avoid conflicts with the future version of the specification.
The PHH files should have a “.phh” filename extension. As for the file naming conventions of standalone hands, we suggest including the last names of the relevant involved players in lowercase characters, with the last names sorted in alphabetical order, each separated by a single dash “-” character, followed by the year the hand took place. For example, the file containing the hand in Figure 1 can be named “dwan-ivey-2009.phh” whereas “arieh-yockey-2019.phh” can be the name of the PHH hand file in Figure 2. As for a bulk of hands for a particular event, the PHH files should all be placed in a folder or nested folders. For an event like 2023 WSOP Event #43: \$50,000 PPC – Day 5 [11], we suggest “wsop/2023/43/5” as the nested folders’ name. For filenames of the files in such folders, we suggest simply using its hand number like “10.phh” for Hand 10 when there is only one table (final table in this case), or using both the hand number and the table number as in “6-51.phh” for Hand 51 in Table 6. These naming conventions are followed in the supplementary PHH files.
X Benchmarking
TABLE VII: The conciseness of each hand history format.
Format # newlines / hand # words / hand # bytes / hand
PHH 9.000 101.634 544.570
PokerStars 34.359 160.519 884.711
Pluribus 1.000 1.000 125.141
Using the average number of newlines, words, and bytes used to represent a hand, we compare the conciseness of the following three formats: the PHH format, the PokerStars hand history format, and the format used in the supplementary of Brown and Sandholm [5] in Table VII. The dataset of hands explored in this analysis is the 10,000 hands played by the poker AI agent Pluribus, all of which were of the no-limit Texas hold ’em variant. The PokerStars hand history version of the dataset was provided by Wang [19] and is used for this comparison. All files exclusively use the ASCII character set.
As expected, the Pluribus’s original format is the most concise of all three representations. This format is followed by the PHH file format, which is noticeably more concise than the PokerStars format, primarily thanks to its non-verbal nature in the recording method.
In order to validate the correctness of the PHH format and our open-source parser implementation, the parsed games were checked for consistency and correctness by comparing the final stacks with the payoffs provided in the dataset of Brown and Sandholm [5].
XI Tooling
from pokerkit import *
# Hand loading
with open("...", "rb") as file:
    hh = HandHistory.load(file)
# Iterate through each action step
for state in hh:
Figure 10: Loading a PHH file in PokerKit [8].
from pokerkit import *
A = Automation
# Game state construction
game = PotLimitOmahaHoldem(
        A.ANTE_POSTING,
        A.BET_COLLECTION,
        A.BLIND_OR_STRADDLE_POSTING,
        A.CARD_BURNING,
        A.HOLE_CARDS_SHOWING_OR_MUCKING,
        A.HAND_KILLING,
        A.CHIPS_PUSHING,
        A.CHIPS_PULLING,
    True,
    0,
    (500, 1000),
    1000,
state = game((1259450.25, 678473.5), 2)
# State progression; Pre-flop
state.deal_hole("Ah3sKsKh")  # Antonius
state.deal_hole("6d9s7d8h")  # Blom
# Blom
state.complete_bet_or_raise_to(3000)
# Antonius
state.complete_bet_or_raise_to(9000)
# Blom
state.complete_bet_or_raise_to(27000)
# Antonius
state.complete_bet_or_raise_to(81000)
state.check_or_call()  # Blom
# Flop
state.deal_board("4s5c2h")
# Antonius
state.complete_bet_or_raise_to(91000)
# Blom
state.complete_bet_or_raise_to(435000)
# Antonius
state.complete_bet_or_raise_to(779000)
state.check_or_call()  # Blom
# Turn & River
state.deal_board("5h")
state.deal_board("9c")
# Creating hand history
hh = HandHistory.from_game_state(
    game, state,
hh.players = [
    "Patrik Antonius", "Viktor Blom",
# Dump hand
with open("...", "wb") as file:
    hh.dump(file)
Figure 11: Dumping to a PHH file in PokerKit [8]. This hand, played between Viktor Blom and Patrik Antonius, holds the record for the largest online poker pot ever played [20].
From PokerKit [8] Version 0.4, the library features PHH file format reader and writer utilities. It offers “load” and “dump” programmatic APIs akin to those provided by Python’s standard libraries such as “json,” and “pickle” [21]. Figures 10 and 11 show sample usages of the PHH file format utilities in PokerKit. The hand history object in Python serves as an iterator of the corresponding poker state which first yields the initial state, followed by the same state after applying each action one-by-one in the “actions” field. From game and state objects that are interacted with programmatically, the hand history object can also be created which can subsequently be saved in the file system.
XII Discussion
The PHH format has the potential to serve as a transformational solution to various computational poker applications involving the storage of poker hands.
XII-A AI Agents
The use of the PHH format can greatly benefit AI research, especially those that use reinforcement learning and ensemble learning techniques. The comprehensive hand histories allow for detailed analyses and can be used for training purposes. Furthermore, using a standardized format to share hand histories played by AI agents can also enhance transparency and enable other researchers to easily validate their findings.
XII-B Online Casinos
Online casinos that offer poker games can store hands played on their platforms in the PHH file format for archival purposes that can track player history and facilitate easier internal auditing processes. The same data (or maybe another version with the sensitive information like other players’ hole cards censored out) can be provided to the players involved in the hand upon request. As an alternative to using persistent files, the developers can map each field into database table columns to store them in a queryable database.
XII-C Poker Tools
The PHH format can become a standard for sharing hands between poker players. One interesting avenue of research is in the creation of an open poker hand database akin to various chess databases which would enable poker enthusiasts to reference different hands. The platform can also serve datasets of poker hands for researchers in the field of computational poker for usage in statistical analyses and AI development.
XIII Conclusion
This paper introduces the PHH file format, designed to concisely describe poker hand history while allowing ease in human usage and machine parser implementations. Our solution fills a long-standing gap in the existing literature on the computer poker field. The specification builds on top of the TOML specification [16], allowing software systems to leverage the rich ecosystem surrounding the TOML file format. By representing different game state information through fields, the hand can be described in a consistent, structured manner. We provide a demonstration of our open-source Python PHH hand parser implementation that utilizes the PokerKit [8] framework as a hand simulator. In the supplementary, we provide over 10,000 PHH files in 11 different diverse variants to validate the format’s flexibility and effectiveness.
The PHH format is suitable to be used in a variety of poker applications in use cases ranging from poker AI development, hand archival, analytical tools, and teaching materials. A promising future direction for research in the field of computational poker is in the development of a poker engine communication protocol akin to the Universal Chess Interface for chess [22] which may utilize numerous portions of the PHH format. The development of such a standard has the potential to accelerate the development of industrial poker AI tools and facilitate their direct comparisons.